OpenStudy (anonymous):

If a thickness of a paper is 9 x10 ^ -3 Cm then how many sheets of paper are in a book of 5 cm thick.

OpenStudy (anonymous):

5cm = 5 x 10^0 cm. So how many times would 9 x 10^-3 go into 5 x 10^0?

OpenStudy (anonymous):

1.8 x 10 ^ -3

OpenStudy (anonymous):

Close, but you're dividing the wrong way.

OpenStudy (anonymous):

so i would have to divide 5 by 9

OpenStudy (anonymous):

which would be 5.5 (bar notation)

OpenStudy (anonymous):

There's a rule for dividing exponents: \[\Large \frac{{{x^n}}}{{{x^m}}} = {x^{n - m}}\] So: \[\Large \begin{array}{l} \frac{{5 \times {{10}^0}}}{{9 \times {{10}^{ - 3}}}} = \frac{5}{9} \times {10^{0 - - 3}} = \frac{5}{9} \times {10^3}\\\\\\ \cong 5.56 \times {10^2} \end{array}\]

OpenStudy (anonymous):

yes but how did you get the 10 ^ 3 to 10^ 2

OpenStudy (anonymous):

5/9 is about 0.556. So 0.556 x 10^3 = 5.56 x 10^2 = 55.6 x 10^1 = 556 x 10^0. When you go down another factor of 10, you move the decimal place one spot to left; if you go up, then one spot to the right. Make sense?

OpenStudy (anonymous):

Oh okay thanks!

OpenStudy (anonymous):

No worries =)